Ryan has the numbers 1–30 listed on individual index cards. If Ryan randomly selects 1 card, what is the probability that it will be a multiple of 3 or a multiple of 11?
Answer by ewatrrr(10682)

 
Hi
1-30
Mulitples of 3: 3,6,9,12,15,18,21,24,27,30 |10 of them
Mulitples of 11: 11,22 |two of them
P(multiple of 3 'or' a multiple of 11? 10/30 + 2/30 = 12/30 = 2/5
Frank has 4 orange paper clips, 6 blue paper clips, 3 red paper clips,
and 5 green paper clips in a box.
If Frank randomly selects 1 paper clip,
what is the probability that he will select 1 red paper clip?
18 paper clips in all
P(1 red clip) = 3/18 = 1/6